Exercise 21-14

Johnson Enterprises uses a computer to handle its sales invoices. Lately, business has been so good that it takes an extra 3 hours per night, plus every third Saturday, to keep up with the volume of sales invoices. Management is considering updating its computer with a faster model that would eliminate all of the overtime processing.
Current Machine New Machine
Original purchase cost $15,200 $25,000
Accumulated depreciation $6,500 _     
Estimated annual operating costs $24,600 $19,900
Remaining useful life 5 years 5 years

If sold now, the current machine would have a salvage value of $12,400. If operated for the remainder of its useful life, the current machine would have zero salvage value. The new machine is expected to have zero salvage value after 5 years.

Should the current machine be replaced? (In the first two columns, enter costs and expenses as positive amounts, and any amounts received as negative amounts. In the third column, enter net income increases as positive amounts and decreases as negative amounts. Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)
Retain
Machine
Replace
Machine
Net Income
Increase
(Decrease)
Operating costs $

$

$

New machine cost

Salvage value (old)

   Total $

$

$

The current machine should be

retained  replaced  

.
Answer #1
Retain Machine Replace Machine Net Income Increase (Decrease)
Operating costs ($24600*5)= $123000 ($19900*5)= $99500 (123000-99500)= 23500
New machine cost 0 25000 -25000
Salvage value (old) 0 -12400 12400
Total $123000 $112100 $10900

The current machine should be Replaced as if the machine will replace the net income will increases.